OpenStudy (anonymous):

anyone can? Find the Taylor series expansion of a function of f(z)=z exp(2z) about z = -1 and find the region of convergence.,

OpenStudy (anonymous):

anyone can? Find the Taylor series expansion of a function of \(f(z)=z e^{2z} \) about z = -1 and find the region of convergence.,

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@Hero @hartnn @RadEn

OpenStudy (anonymous):

I think I can get this as soon as I review on Taylor series

OpenStudy (anonymous):

to find the region of convergence look here http://faculty.eicc.edu/bwood/ma155supplemental/supplemental26.htm

OpenStudy (anonymous):

it shows you how

OpenStudy (anonymous):

I think I need to see someone else do this, I've forgotten a lot since last year.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

should we assume z as a general variable, or a complex variable?

OpenStudy (amistre64):

\[e^u=\sum_0\frac{1}{n!}u^n\] \[ue^u=\sum_0\frac{1}{n!}u^{n+1}\] \[ue^u=\sum_0\frac{1}{n!}(u+1)^{n+1}\] maybe

OpenStudy (amistre64):

region of convergence i would assume is similar to radius or interval of convergence, but in a disc shape

OpenStudy (anonymous):

and then ???

OpenStudy (amistre64):

not too sure, im not that confident in imaginary calculus .... from what i seened online, the rules are the same.

OpenStudy (anonymous):

can you help me for this ??? @UnkleRhaukus @nubeer @electrokid @jhonyy9 @mathslover